# RNA-FrameFlow: Flow Matching for de novo 3D RNA Backbone Design

### v1

@@ B.2 RhoFold Length Bias
We investigate the performance of RhoFold on the training dataset used for our generative model. Figure 11 shows that RhoFold has a sequence length bias where it predicts good structures with low RMSDs for specific sequence lengths (like 70, 100, and 120) while predicting poor structures for other lengths with larger RMSDs. The performance across lengths is disparate and may influence what is considered ‘valid’ in our unconditional generation benchmarks.
Figure 11: RhoFold length bias. RhoFold has a strong bias for certain sequence lengths over others. This affects its efficacy when used to compute the 3D self-consistency of generated backbones. The blue dotted line represents the median RMSD of RhoFold predictions to the samples. To minimize the influence of this length bias, we use TM-score for self-consistency because it does not penalize flexible regions as much as RMSD.
Appendix C Ablation Study
C.1 Composition of Backbone Coordinate Loss
@@
We also analyze how changing the composition of atoms considered in the inter-atom losses affects performance. We increase the number of atoms being supervised in the $\mathcal{L}_{\text{bb}}$ loss described above. Aside from the frame comprising $C3^{\prime}$, $C4^{\prime}$, and $O4^{\prime}$, we try two settings with 3 and 7 additional non-frame atoms included in the loss. For the 3 non-frame atoms, we additionally choose $C1^{\prime}$, $P$, and $O3^{\prime}$, and for the 7 non-frame atoms, we choose a superset $C1^{\prime}$, $P$, $O3^{\prime}$, $C5^{\prime}$, $OP1$, $OP2$, and $N1/N9$. We posit the additional supervision may increase the local structural realism, which may further improve validity, as shown in Table 5.
@@
Table 5: Ablating composition of backbone loss $\mathcal{L}_{\text{bb}}$. Supervising more non-frame atoms improves validity but worsens diversity and novelty. Best per-column result is bolded.
@@
C.2 Composition of Auxiliary Loss
We ablate the inclusion of different auxiliary loss terms that guide our $SE(3)$ flow matching setup; results are in Table 6. Although, there is an increase in EMD for bond distances as we remove distance-based losses like backbone coordinate loss $\mathcal{L}_{\text{bb}}$ and all-to-all pairwise distance loss ($\mathcal{L}_{\text{dist}}$). However, we also observe the model still learns realistic distributions despite removing different loss terms, indicating that each loss makes up for the absence of the other. Moreover, the best model still uses all losses with any removal causing a drop in validity.
Further inspecting the samples from the models without each loss term reveals structural deformities at the all-atom level. Figure 12 shows such artifacts resulting from not enforcing geometric constraints through explicit losses.
@@ C.2 Composition of Auxiliary Loss
Figure 12: Not including auxiliary losses causes structural deformities in generated RNAs. (A) RNA backbone from our baseline model with expected adherence to bonding between nucleotides. (B) Not including the rotation loss $\mathcal{L}_{SO(3)}$ causes nucleotides to have random orientations, preventing them from connecting contiguously. (C) Not including the backbone atom loss $\mathcal{L}_{\text{bb}}$ causes intra-residue atoms to be placed too close to one another resulting in bonds that should not exist. (D) Not including the all-to-all pairwise distance loss $\mathcal{L}_{\text{dist}}$ causes deformations and fusing between adjacent frames, and unrealistic nucleotide placements, especially along helices and loops.
Appendix D Additional Results
D.1 Evaluation of MMDiff Samples
@@
D.2 Evaluation of Data Preparation Strategies

### v3

@@ A.3 RhoFold Length Bias
We investigate the performance of RhoFold on a representative subset of the training dataset used to train RNA-FrameFlow. Figure 12 shows that RhoFold has a sequence length bias where it predicts accurate structures with low RMSDs (to the ground truth) for specific sequence lengths (like 70, 100, and 120) while predicting poor structures for other lengths. The performance across lengths is disparate and may influence what is considered valid in our unconditional generation benchmarks. This affects its efficacy when used in a self-consistency pipeline with the RMSD metric. To minimize the influence of this length bias, we use TM-score for self-consistency because it does not penalize flexible regions as much as RMSD.
Figure 12: RhoFold length bias. The blue dotted line represents the median RMSD of RhoFold predictions to the RNAsolo samples.
A.4 Imputing Non-frame Atoms from Torsion Angles
Here, we describe how we autoregressively impute the remaining non-frame atoms using 8 torsion angles $\Phi=\{\phi_{1}\rightarrow\phi_{8}\}$. For a nucleotide $n$ along the generated RNA backbone, we assume we have its final frame $T^{(n)}=(r^{(n)},x^{(n)})$ obtained from the denoiser’s output after $N_{T}$ diffusion timesteps. Going by our choice of frame $\{C3^{\prime},C4^{\prime},O4^{\prime}\}$, we place non-frame atoms in the following order: $C2^{\prime},C1^{\prime},N1/N9,O3^{\prime},O5^{\prime},P,OP1,OP2$, each corresponding to its respective $\phi_{i}\in\Phi$ as shown in Figure 1.
Referring to the figure on the right, suppose we have three atoms A, B, and C with coordinates $(x_{A},y_{A},z_{A}),(x_{B},y_{B},z_{B}),(x_{C},y_{C},z_{C})$. They are connected by bonds AB and BC denoted by vectors $\overrightarrow{AB}=B-A$ and $\overrightarrow{BC}=C-B$ with lengths $r_{AB}=|\overrightarrow{AB}|$ and $r_{BC}=|\overrightarrow{BC}|$ respectively. To rotate $BC$ around $AB$ by some angle $\phi$, we perform the following procedure:
  1. 1.
    Compute unit vector $\vec{u}_{AB}=\frac{\overrightarrow{AB}}{r_{AB}}$ along bond AB by normalizing $\overrightarrow{AB}$.
  2. 2.
    Compute a vector perpendicular to $\vec{u}_{AB}$ by choosing a random normal vector $\vec{n}$ (like $[1,0,0]^{T}$ or $[0,1,0]^{T}$) and taking their cross product to get $\vec{u}_{1}=\vec{u}_{AB}\times\vec{n}$.
  3. 3.
    Compute the unit vector $\vec{u}_{\text{rot}}$ rotated by $\phi$ around $AB$ using Rodrigues’ rotation formula:
    $\vec{u}_{\text{rot}}=\cos{(\phi)}\cdot\vec{u}_{1}+\sin{(\phi)}\cdot(\vec{u}_{AB}\times\vec{u}_{1})+(1-\cos{(\phi)})(\vec{u}_{AB}\cdot\vec{u}_{1})\cdot\vec{u}_{AB}\leavevmode\nobreak\ .$
  4. 4.
    Compute the coordinates of atom $C$ as follows:
    $(x_{C},y_{C},z_{C})=(x_{B},y_{B},z_{B})+r_{BC}\cdot\vec{u}_{\text{rot}}\leavevmode\nobreak\ .$
We use predetermined bond lengths between atoms in the idealized geometry of the Adenine (A) nucleotide from OpenComplex (Jingcheng et al., 2022) in the same way Yim et al. (2023b; a) use Alanine for generated protein backbones. We use the following atom triplets and predicted torsion angles to build the all-atom nucleotide, starting from the ribose sugar ring towards the $5^{\prime}$ end:
Fixed bond Non-frame atom Torsion angle
$C4^{\prime}-C3^{\prime}$ $C2^{\prime}$ $\phi_{1}$
$C4^{\prime}-O4^{\prime}$ $C1^{\prime}$ $\phi_{2}$
$O4^{\prime}-C1^{\prime}$ $N9$ (or $N1$) $\phi_{3}$
$C4^{\prime}-C3^{\prime}$ $O3^{\prime}$ $\phi_{4}$
$C4^{\prime}-C5^{\prime}$ $O5^{\prime}$ $\phi_{5}$
$C5^{\prime}-O5^{\prime}$ $P$ $\phi_{6}$
$O5^{\prime}-P$ $OP1$ $\phi_{7}$
$O5^{\prime}-P$ $OP2$ $\phi_{8}$
Appendix B Ablations
B.1 Composition of Backbone Coordinate Loss
@@
We also analyze how changing the composition of atoms in the inter-atom losses affects performance. We increase the number of atoms being supervised in the $\mathcal{L}_{\text{bb}}$ loss described above. Aside from the frame comprising $C3^{\prime}$, $C4^{\prime}$, and $O4^{\prime}$, we try two settings with 3 and 7 additional non-frame atoms included in the loss. For the 3 non-frame atoms, we additionally choose $C1^{\prime}$, $P$, and $O3^{\prime}$, and for the 7 non-frame atoms, we choose a superset $C1^{\prime}$, $P$, $O3^{\prime}$, $C5^{\prime}$, $OP1$, $OP2$, and $N1/N9$. We posit the additional supervision may increase the local structural realism, which may further improve validity, as shown in Table 5.
@@
Table 5: Ablating composition of backbone loss $\mathcal{L}_{\text{bb}}$. Supervising more non-frame atoms improves validity but worsens diversity and novelty. Best result per column is highlighted.
B.2 Composition of Auxiliary Loss
We ablate the inclusion of different auxiliary loss terms that guide our $SE(3)$ flow matching setup; results are in Table 6. Although, there is an increase in EMD for bond distances as we remove distance-based losses like backbone coordinate loss $\mathcal{L}_{\text{bb}}$ and all-to-all pairwise distance loss ($\mathcal{L}_{\text{dist}}$). However, we also observe the model still learns realistic distributions despite removing different loss terms, indicating that each loss makes up for the absence of the other. Moreover, the best model still uses all losses with any removal causing a drop in validity. Further inspecting the samples from the models without each loss term reveals structural deformities at the all-atom level. Figure 13 shows such artifacts resulting from not enforcing geometric constraints through explicit losses.
@@ B.2 Composition of Auxiliary Loss
Figure 13: Not including auxiliary losses causes structural deformities in generated RNAs. (A) RNA backbone from our baseline model with expected adherence to bonding between nucleotides. (B) Not including the rotation loss $\mathcal{L}_{SO(3)}$ causes nucleotides to have random orientations, preventing them from connecting contiguously. (C) Not including the backbone atom loss $\mathcal{L}_{\text{bb}}$ places intra-residue atoms too close to one another resulting in bonds that should not exist. (D) Not including the all-to-all pairwise distance loss $\mathcal{L}_{\text{dist}}$ causes adjacent frames to fuse and loses contiguity, especially along helices and loops.
B.3 Choice of Forward-folding Model
In our work, we rely on RhoFold (Shen et al., 2022) to forward fold the inverse-folded sequences from gRNAde. Here, we reperform our evaluation from Section 4.1 with Chai-1 (Boitreaud et al., 2024), a recent open-source structure prediction model with results similar to AlphaFold2, replacing RhoFold in the self-consistency pipeline in Figure 2. We do not use MSAs for Chai-1.
We do not observe any major differences in self-consistency between the two forward-folding models. For RNA-FrameFlow with RhoFold, we report a validity of 41.0% while RNA-FrameFlow with Chai-1 gives a validity of 39.5%. Recent benchmarks (Tarafder et al., 2024) also observe that existing RNA structure prediction tools like RhoFold, RF2NA (Baek et al., 2022b), and trRosettaRNA (Wang et al., 2023) perform similarly due to similarities in their architectures and training data. We compare the scTM distribution among valid samples between RhoFold and Chai-1 predictions in single-sequence mode in Figure 14.
Figure 14: Correlation between RhoFold and Chai-1 scTM scores. Horizontal and vertical dotted lines indicate the average scTM score from each method among valid samples.
Appendix C Additional Results
C.1 Evaluation of MMDiff Samples
@@
C.2 Evaluation of Data Preparation Strategies
@@
C.3 Comprehensive local evaluation of angular distributions
Following the empirical structural analysis of RNA by Gelbin et al. (1996), we compare local bond angle distributions among triplets of atoms from the generated backbones. We sample 50 all-atom backbones for each sequence length in $[70,90,110,130,150]$, sieve out the valid samples, and extract relevant bond angles. As shown in Figure 19, we observe that RNA-FrameFlow can retrieve angular distributions between distant and nearby atoms in the nucleotides, providing preliminary evidence that modern protein design models are sufficiently expressive to model RNA tertiary structure.
Figure 19: Bond angle distributions between triplets of atoms. We select these atomic triplets from the empirical study of RNA’s 3D geometry by Gelbin et al. (1996).
C.4 Measuring All-atom Steric Clashes
We compare the all-atom-level steric clashes between filtered RNAsolo samples used for training and the generated backbones from RNA-FrameFlow. We say two unbonded atoms $i,j$ clash if the distance between them $r_{ij}$ is within a threshold $d_{\text{steric}}$:
$d_{\text{steric}}$ $=\text{v}_{i}+\text{v}_{j}-0.6$ (11)
$\mathbb{I}_{ij}$ =$=\begin{cases}1&r_{ij}\leq d_{\text{steric}}\\ 0&\text{otherwise}\end{cases}$ (12)
# clashes $=\sum_{i,j}\mathbb{I}_{ij}\ .$ (13)
Here, $\text{v}_{i},\text{v}_{j}\in\mathbb{R}$ are the Van der Waals (VdW) radius of the atoms $i,j$ in Angstrom. Based on its identity, each atom has its own VdW radius which we factor into our computation. We leave a generous tolerance of 0.6 Å  (corresponding to half the Hydrogen atom’s VdW radius of 1.20 Å) to account for random deviations in atomic placements. We ignore Phosphodiester and Glycosidic bonds when computing clashes because the covalent radius is smaller than the VdW radius. As nucleotides are constructed using idealized bonds, there may be fewer inter-nucleotide clashes, resulting in fewer clashes for RNA-FrameFlow backbones.
In Figure 20, we compare the steric clashes across sequence length bins. We observe that RNA-FrameFlow generates backbones that have a similar distribution of inter-atom steric clashes as samples from RNAsolo. We also include validity for each sequence length bucket. We see that samples from certain sequence lengths (like 70, 80, 120) contain relatively fewer steric clashes across samples within that length bucket since they are over-represented in RNAsolo. This means RNA-FrameFlow might be better at recapitulating atomic positions for such lengths than others. The steric clashes are normalized by the number of heavy atoms in the molecules, giving us steric clashes per 100 atoms. For the RNAsolo samples, we see $10.03\pm 1.52$ clashes per 100 atoms while our generated backbones have $25.55\pm 5.43$ clashes per 100 atoms.
Figure 20: All-atom steric clashes by sequence length. We observe a similar number of steric clashes between training and generated backbones across sequence lengths. We include the (validity) for generated samples from each sequence length below the length labels along the horizontal axis.